If demand function is given as the following:

 

Qz = 230 -2.75 Pz + 0.5 I + 1.2 Pm + 0.6A

Where Qz is quantity of Good z sold, Pz is price of Good z per unit, I is per capita income, Pm is price of competitor and A is the amount of advertising spent.

Current values:  Pz= RM 55 I= RM 9000    Pm= RM 50    A =RM 12,000.

a)Determine the equation for demand, total revenue and marginal revenue curve.

b)Determine the total revenue maximizing price and quantity.


1
Expert's answer
2021-04-28T20:14:38-0400
"Q_Z=230-2.75p_z+0.5\times9000+1.2\times50+0.6\times12000"

"Q_Z=11990-2.75p_Z"

"TR=(11990-2.75p_Z)p_Z"

"MR=11990-5.5p_z"

"p_Z=2180"

"Q_Z=5995"